Calculate the molarity of a solution containing 2.40g of NaCl in 40.0 mL of solution. (MM NaCl = 58.45g/mol).


1
Expert's answer
2021-02-23T04:36:57-0500

moles of NaCl = 2.40/58.45 = 0.041 moles

Molarity = moles of solute / volume of solution

Volume = 40ml = 0.040L

M = 0.041/0.040 = 1.025M